Completing the square.
Please walk me through each step. the text says the answers are x=10, x=-2
x^2 - 8x - 20 = 0
thank you

Answer by ewatrrr(24785)


Hi
x^2 - 8x - 20 = 0 |Completing the square
(x-4)^2 -16 - 20 = 0 |Note:(x-4)^2 -16 = (x^2-8x+16) -16 = x^2-8x
(x-4)^2 = 36
(x-4) = ± 6
x-4 = 6
x = 10
x-4 = -6
x = -2